# GTPO: Trajectory-Based Policy Optimization in Large Language Models

### v3

@@ Abstract
Policy-based optimizations are widely adopted today for the training and alignment of language models, where one of the most recent and effective approaches is Group-relative Policy Optimization (GRPO). In this paper, we reveal and analyze two major limitations of GRPO: (i) tokens frequently appear in completions with both positive and negative rewards, leading to conflicting gradient updates that can reduce their output probability, even though can be essential for maintaining proper structure; (ii) negatively rewarded completions may penalize confident responses and shift model decisions toward unlikely tokens, progressively flattening the output distribution and degrading learning. To address these issues and provide a more stable and effective policy optimization strategy, we introduce GTPO (Group-relative Trajectory-based Policy Optimization), which identifies conflict tokens, tokens appearing in the same position across completions with opposite rewards and protects them by skipping negative updates, while amplifying positive ones. To further prevent policy collapse, GTPO filters out completions whose entropy exceeds a provable threshold. Unlike GRPO, GTPO does not rely on KL-divergence regularization, eliminating the need for a reference model during training, while still ensuring greater training stability and improved performance, validated through multiple experiments on GSM8K, MATH and AIME 2024 benchmarks. The code is available on Github at this link [^1] http://bit.ly/4oV8c49.
@@ 1 Introduction
Recent advancements in the training and alignment of large language model (LLM) have led to the adoption of policy-based optimization techniques to encourage LLMs on matching desired human preferences. Prominent approaches such as DPO (Rafailov et al. 2023), RLHF (Ouyang et al. 2022), and RFT (Yuan et al. 2023) were among the first to incorporate human feedback to guide model responses during training, acting as Reinforcement Learning (RL) techniques (Shao et al. 2024). More recently, a significant advancement in RL was introduced with Group-Relative Policy Optimization (GRPO) (Shao et al. 2024), which represents a variant of PPO (Schulman et al. 2017) that eliminates the need for a specific critic model by estimating the baseline directly from grouped reward values. In GRPO, the LLM acts as a policy that generates step-by-step reasoning and receives deterministic rewards based on the aggregated correctness and formatting of multiple tentative answers (referred to as completions) inferred for each question.
@@
Despite the improvements introduced by GRPO, there are two key issues identified and analyzed in this work: (i) undesired gradient conflicts affect potentially corrects tokens shared across multiple completions of the same group that receive both positive and negative advantages, and (ii) a policy collapse phenomenon, defined as a degradation in LLM performance after a certain number of training steps, in which negatively rewarded completions destabilize the training process. More specifically, we observe that the first issue primarily impacts formatting tokens, which are essential to ensure proper answer structure and style.
@@
To address these issues, this paper proposes Group-relative Trajectory-based Policy Optimization (GTPO), which treats the sequence of generated tokens (i.e., the completion) as a trajectory of decisions taken by the LLM policy. The core idea behind GTPO is to prevent undesired divergence among trajectories within the same group, thus keeping stability, while enhancing the rewards. To do so, GTPO applies a conflict-aware gradient correction mechanism that mitigates gradient conflicts on shared tokens, particularly in the initial and final parts of completions, thus preserving formatting consistency across trajectories. Furthermore, we demonstrate that while KL divergence of GRPO often reacts too slowly in preventing policy collapse, monitoring the entropy in LLM outputs offers a clearer signal of policy instability. Building on this, entropy-based regularization terms are applied in GTPO to control the exploration of trajectories in the same group.
These two components introduced by GTPO prevents the penalization of important formatting tokens and mitigates policy collapse, improving both the structure and accuracy of generated completions. Notably, since GTPO no longer relies on KL divergence, it does not require a reference model during training, making the process more lightweight and faster. In summary, the contributions of the work are:
@@
  • We identify and analyze two critical issues in GRPO, such as gradient conflicts on tokens shared between completions of the same groups; and a policy collapse phenomenon, highlighting the limitations of the KL term;
@@
  • We introduce GTPO, a new policy-based optimization that addresses the previous issues via conflict-aware gradient corrections and entropy-based regularizations;
@@
  • We conduct extensive experiments and ablations studies to validate the effectiveness of GTPO, demonstrating more stable training and improved performance on both in-distribution and out-of-distribution benchmarks, including GSM8K, MATH, and AIME2024, using LLaMA-8B and Qwen 2.5-3B.
The remainder of the paper is organized as follows: we first discuss related work and introduce important preliminaries. We then analyze and discuss issues identified in GRPO, followed by the description of GTPO. Finally, the paper presents the experimental results and states conclusions.
@@ 2 Related Work
Reinforcement Learning in LLMs. RL has been widely adopted in decision-making tasks (Mnih et al. 2016, 2015; Berner et al. 2019), and nowadays is increasingly applied to the alignment and fine-tuning of LLMs. A prominent approach is RL from Human Feedback (RLHF), introduced in InstructGPT (Ouyang et al. 2022), and further developed by Anthropic (Bai et al. 2022). RLHF has become central to the training pipelines of state-of-the-art LLMs such as Claude 3 (Anthropic 2024), Gemini (Anil et al. 2023), and GPT-4 (OpenAI 2023). It typically includes supervised fine-tuning, a reward model, and the adoption of Proximal Policy Optimization (PPO) (Schulman et al. 2017). In this settings, PPO improves training stability by constraining updates through a clipped surrogate objective, offering a more practical alternative to Trust Region Policy Optimization (TRPO) (Schulman et al. 2015). However, it remains sensitive to reward scaling and can suffer from training instability (Wang et al. 2019; Garg et al. 2021; Moalla et al. 2024), requiring multiple refinements (Huang et al. 2022). Therefore, multiple variations have been proposed over the years, such as TRGPPO (Wang et al. 2019), alphaPPO (Xu et al. 2023), and PPO-ALR (Jia et al. 2024).
Advancements and Limitations in GRPO. To overcome the need for a critic model, Deepseek introduced GRPO (Shao et al. 2024; Guo et al. 2025), an approach that, given a question, compares multiple responses (completions) to derive relative rewards. GRPO demonstrates state-of-the-art performance on math benchmarks and achieves human-like alignment without relying on explicit manual feedback or critic networks (Li et al. 2025). Despite its promise, potential limitations of GRPO have been recently emerged, as bias effects (He, Fried, and Welleck 2025), gradient imbalance (Yang et al. 2025b), which lead to undertraining of rare yet informative tokens (Liu et al. 2025), and degradations (or even collapse) of model performance like in PPO (Dohare, Lan, and Mahmood 2023).
Building upon previous analyses of GRPO training behaviors, we deepen the study of token-level updates across completions of the same group, revealing conflicts in shared tokens. Furthermore, we extend the understanding of policy collapse, showing that KL divergence is limited in addressing this issue, whereas entropy-based analysis provides clearer signals (Cui et al. 2025). These insights motivate the design of GTPO, which effectively improves stability and performance during both training and evaluation.
@@ 3 Preliminaries
In GRPO, the LLM, acting as a policy, generates during training multiple completions (responses) per prompt $q$, denoted as $\{o_{1},o_{2},\ldots,o_{G}\}$, where $G$ is the number of completions generated, and computing rewards relatively to each rather than absolutely. Each output is structured with special tags for reasoning and answer segments and allows the reward to be calculated as an aggregate of formatting and factual correctness. The goal is to maximize the objective:
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{{o}_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\bar{\mathcal{C}}_{i}-\beta\cdot D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})\right]$ (1)
Here, the expectation is taken over prompts $q\sim\mathbb{P}(Q)$ and completions $\{o_{i}\}\sim\pi_{\theta_{\text{old}}}$, where $\pi_{\theta_{\text{old}}}$ is the behavior policy. Each completion $o_{i}$ has length $|o_{i}|$. The term $\bar{\mathcal{C}}_{i}$ denotes the average clipped advantage over its tokens: $\bar{\mathcal{C}}_{i}=\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\mathcal{C}_{i,t}.$ The clipped advantage for the token at position $t$ in the completion $o_{i}$, is defined as: $\mathcal{C}_{i,t}=\min\left(\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}({o}_{i,t}\mid s_{i,t})}\cdot\hat{A}_{i},\ \text{clip}\right),$ where $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\mathrm{std}(R)}$ is the normalized scalar advantage assigned to the $i$-th completion, and $s_{i,t}=(q,o_{i,<t})$ denotes the sequence of previously generated tokens up to position $t$. Specifically $R_{i}$ represents the reward for completion $i$, which is a sum of sub-rewards based on the correctness of the answer and its formatting style (Formatting Reward). The term $\pi_{\theta}(o_{i,t}|s_{i,t})=\mathrm{softmax}(f_{\theta})\in\mathbb{R}^{V},$ represents the probability distribution over the output logits $f_{\theta}$, where $f^{j}_{\theta}$ is the logit at position $j$, given the context $s_{i,t}$, and $V$ is the vocabulary size. Finally, the KL divergence term in Eq.(1) penalizes deviation from a reference policy $\pi_{\mathrm{ref}}$, scaled by a factor $\beta$. Due to computational cost and training time, one iteration is adopted in Eq. 1 (Simoni et al. 2025), where $\pi_{\theta}=\pi_{\theta_{\text{old}}}$, making the likelihood ratio 1 and the clipping unnecessary, i.e., $\mathcal{C}_{i,t}=\hat{A}_{i}$. The gradient of this simplified objective is:
@@
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{A}_{i}}{|{o}_{i}|}\sum_{t=1}^{|{o}_{i}|}g_{i,t}-\beta\cdot\nabla_{\theta}\mathrm{D}_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$ (2)
where $g_{i,t}:=\nabla_{\theta}\log\pi_{\theta}({o}_{i,t}|s_{i,t})$ denotes the token-level policy gradient, which can be also expressed as:
$g_{i,t}=\Big{(}\nabla_{\theta}f_{\theta}^{j}(1-\pi_{\theta}^{j})-\sum_{k\neq j}\pi_{\theta}^{k}\nabla_{\theta}f_{\theta}^{k}\Big{)},$
with $j$ is the index of the logit with the highest probability $\pi_{\theta}$. Intuitively from Eq.(2), a positive advantage ($\hat{A}_{i}>0$) boosts the logit of the selected token proportionally to $1-\pi^{j}$, while reducing the logits of all others. Converselly, a negative advantage inverts this behavior, penalizing the token and increasing alternatives. Full derivations are in Append.
@@ 4 GRPO Issues
This section highlights two major limitations of GRPO: Token-level penalization and policy collapse.
@@ 4.1 Token-level Penalization
We show that GRPO tends to have negative effects on the updates of certain tokens that are essential for maintaining the structure and interpretability of completions.
@@
Prefix Tokens Penalization. Given a prompt $q$, the model generates a set of completions $\{o_{1},o_{2},\ldots,o_{G}\}$, which may begin with the same sequence of tokens and then diverge at a specific point, a token that acts as a crossroads. This behavior is typical in instruction-tuned models, where early tokens tend to be similar across completions (Wang et al. 2024). However, not all completions in the group $G$ necessarily share the same prefix. In practice, we often observe multiple distinct prefixes, each shared by a subset of the completions. To model this, we partition the $G$ completions into $K$ disjoint groups $\mathcal{G}_{1},\ldots,\mathcal{G}_{K}$ so that all completions in a group $\mathcal{G}_{k}$ have a common prefix $S_{\text{pfx}}^{(k)}$. If all $G$ completions share the same prefix, then $K=1$. Tokens in $S_{\text{pfx}}^{(k)}$ are affected by the combined advantages of all completions in the group, while other tokens only depend on the advantage of their own completions. Under this structure, we rewrite the GRPO gradient (Eq. 2) by summing the contributions of all groups, where each group contributes a prefix term and a set of individual terms. The gradient, excluding the KL term, becomes:
@@
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{k=1}^{K}\Bigg{[}\underbrace{\sum_{i\in\mathcal{G}_{k}}\frac{\hat{\mathcal{A}}_{i}}{|{o}_{i}|}\sum_{t\in S_{\text{pfx}}^{(k)}}g_{i,t}}_{\Delta_{\text{pfx}}^{(k)}}+\sum_{i\in\mathcal{G}_{k}}\frac{\hat{\mathcal{A}}_{i}}{|{o}_{i}|}\sum_{j\notin S_{\text{pfx}}^{(k)}}g_{i,j}\Bigg{]}$
The term $\Delta_{\text{pfx}}^{(k)}$ represents the gradient update associated with the shared prefix of group $\mathcal{G}_{k}$. If a completion does not share any prefix with others, then $\mathcal{G}_{k}$ contains only that single completion, resulting in $S_{\text{pfx}}^{(k)}=\emptyset$ and $\Delta_{\text{pfx}}^{(k)}=0$.
@@
Since the log-probability terms $g_{i,t}$ are identical and so constant across completions in the group $S_{\text{pfx}}^{(k)}$, the gradient component $\Delta_{\text{pfx}}$ is primarily driven by the sum of normalized advantages $\sum_{i\in\mathcal{G}_{k}}\hat{\mathcal{A}}_{i}/|{o}_{i}|$. This implies that the update to a prefix mainly depends on the relative balance of advantages across all completions in group $k$. For instance, the sum can be negative, when correct completions (with positive advantages) are generally longer than incorrect ones. In such cases, the denominator $|{o}_{i}|$ for the correct completions becomes larger, reducing their contribution to the overall gradient. As a result, the model may be penalized for generating desirable and beneficial prefix tokens. This introduces a bias against shared initial tokens, such as formatting tokens or reasoning tags (<reasoning>), which are essential for the structure and correctness of completions.
Dependencies From Completion Advantage. GRPO can also induce undesirable penalization on tokens that appear after the prefix. Certain tokens, in particular formatting ones, may occur in multiple completions at varying positions and within different contexts, some correct, others incorrect, making them more susceptible to inconsistent and potentially harmful updates.
@@
This induces potential issues in the update. For instance, if a token $\tau$ frequently appears in completions with negative advantage, its probability may be reduced, even if the token is syntactically correct and required. Additionally, further issues can arise when a completion only partially follows the expected format. For example, if a completion closes a first part correctly with a formatting token $\tau$ </reasoning> but then omits <answer>, $\tau$ may receive a low or even negative reward. This, in turn, penalizes </reasoning>, despite it being a correct token in the completion.
This occurs because the update for each token primarily depends on the advantage assigned to the entire completions in which it appears, without directly considering whether the token’s individual contribution was beneficial. We acknowledge that this phenomenon is particularly impactful in the case of formatting tokens, as illustrated in the previous examples, and in the case of shared final tokens that form a common suffix among completions.

### v4

# GTPO: Stabilizing Group Relative Policy Optimization via Gradient and Entropy Control

@@ Abstract
Group Relative Policy Optimization (GRPO) is a promising policy-based approach for Large Language Model alignment, yet its performance is often limited by training instability and suboptimal convergence. In this paper, we identify and analyze two main GRPO issues: (i) the token-level penalization, where valuable tokens shared across different responses receive contradictory feedback signals, leading to conflicting gradient updates that can reduce their likelihood; and (ii) the policy collapse, where negatively rewarded completions may penalize confident responses and shift model decisions toward unlikely tokens, destabilizing training process. To address these issues we introduce GTPO (Group-relative Trajectory-based Policy Optimization), which prevents conflicting gradients on valuable tokens by skipping negative updates while amplifying positive ones and filters out completions whose entropy exceeds a provable threshold, to prevent policy collapse. Unlike GRPO, GTPO does not rely on KL-divergence regularization, eliminating the need for a reference model during training, while still ensuring greater training stability and improved performance, as validated through multiple experiments on GSM8K, MATH, AIME 2024, AIME 2025 and AMC 2023. The Github code is available here [^1] https://github.com/winstonsmith1897/GTPO..
@@ 1 Introduction
Recent advances in the training of large language models (LLMs) have led to the use of policy-based optimization methods, where the model is treated as a policy that is updated to better match human preferences. Methods such as DPO DPO, RLHF rlhf, and RFT RFT follow this idea: they collect human or heuristic feedback on model outputs and use it as a reward signal to increase the likelihood of preferred or desired responses. Group Relative Policy Optimization (GRPO) GRPO is a more recent RL-based method that builds on PPO ppo but removes the need for a separate critic network, i.e., the value function estimator. GRPO compares multiple completions for the same question, each consisting of natural-language reasoning followed by the final answer, and computes advantages from their relative rewards based on completion correctness and formatting quality. The rewards are computed using RL with Verifiable Rewards (RLVR) GRPO, ensuring that they are deterministic and verifiable.
@@
Despite the improvements introduced by GRPO, there are two key issues identified and analyzed in this work: (i) undesired gradient conflicts affect potentially correct tokens shared across multiple completions of the same group that receive both positive and negative advantages, and (ii) a policy collapse phenomenon, defined as a degradation in policy performance during training, in which negatively rewarded completions may penalize confident responses and flatten tokens likelihood distribution, shifting model decisions toward unlikely tokens.
@@
To address these issues, we propose Group-relative Trajectory-based Policy Optimization (GTPO), which treats the sequence of generated tokens (i.e., the completion) as a trajectory of decisions taken by the LLM policy. The core idea behind GTPO is to control gradient conflict and entropy explosion while keeping the average output entropy from becoming too low, so that exploration is not prematurely suppressed. To do so, GTPO applies a conflict-aware gradient correction mechanism that mitigates gradient conflicts on shared tokens, particularly in the initial and final parts of completions, thus preserving consistency across trajectories. Furthermore, we show that while the Kullback–Leibler (KL) divergence of GRPO often reacts too slowly in preventing policy collapse, monitoring the average entropy of the output tokens distribution offers a clearer signal of training instability. Building on this, entropy-based regularization terms are applied in GTPO to control the exploration of trajectories in the same group, also filtering out completions (regardless of their corresponding advantage) whose entropy exceeds a provable threshold. In this way, GTPO prevents the penalization of important tokens and mitigates policy collapse, improving both the formatting and accuracy of generated completions. Moreover, GTPO removes the KL divergence term used in vanilla GRPO GRPO, avoiding the reference model and making training faster. In summary, the contributions of the work are:
@@
  • We identify two critical issues in GRPO: (i) gradient conflicts on tokens shared across completions within the same group, and (ii) a policy collapse phenomenon that exposes the limitations of the KL term.
@@
  • We introduce GTPO, a novel policy-based optimization that addresses the previous issues via conflict-aware gradient corrections and entropy control.
@@
  • We conduct extensive experiments and ablations studies on mathematical reasoning task to validate the effectiveness of GTPO against GRPO and Supervised fine-tuning (SFT), using LLaMA 8B llama and Qwen 2.5 (3B) qwen. We demonstrate more stable training and improved performance on both in-distribution, GSM8K GSM8K and MATH MATH, and out-of-distribution benchmarks, AIME 2024 aime2024, AIME 2025 aime2025 and AMC 2023 amc2023. In addition, ablation studies confirm the importance of each term of GTPO.
The paper is structured as follows: Section 2 reviews related work, and Section 3 introduces key preliminaries. Section 4 analyzes GRPO’s limitations, while Section 5 presents our GTPO approach. Section 6 reports experimental results, and Section 7 concludes the paper.
@@ 2 Related Work
Reinforcement Learning in LLMs. RL has been widely adopted in decision-making tasks google-rl; rlhf-notLLM; dota-2, and nowadays is increasingly applied to the alignment and fine-tuning of LLMs. A prominent approach is RL from Human Feedback (RLHF), introduced in InstructGPT rlhf, and further developed by Anthropic antropic-rlhf. RLHF has become central to the training pipelines of state-of-the-art LLMs such as Claude 3 anthropic2024claude3, Gemini gemini, and GPT-4 GPT4. It typically includes supervised fine-tuning, a reward model, and the adoption of Proximal Policy Optimization (PPO) ppo. Here, PPO improves training stability by constraining updates through a clipped surrogate objective, offering a more practical alternative to Trust Region Policy Optimization (TRPO) TRPO. However, it remains sensitive to reward scaling and can suffer from training instability trgppo; PPOissues2; PPOissues, requiring multiple refinements ppo37implementations. Therefore, multiple variations have been proposed over the years, such as TRGPPO trgppo, alphaPPO alphaPPO, and PPO-ALR PPO-ALR.
Advancements and Limitations in GRPO. To overcome the need for a critic model, Deepseek introduced GRPO GRPO; GRPO-r1, an approach that, given a question, compares multiple responses (completions) to derive relative rewards. GRPO demonstrates state-of-the-art performance on math benchmarks and achieves human-like alignment without relying on explicit manual feedback or critic networks grpo-alignment-human. Despite its promise, potential limitations of GRPO have been recently emerged, as bias effects overfit-grpo, gradient imbalance low-token-grpo-issue, which lead to undertraining of rare yet informative tokens liu2025understanding, and degradations (or even collapse) of model performance like in PPO dohare2023overcoming. Recently, Group Sequence Policy Optimization (GSPO) zheng2025group has been proposed as a GRPO variant that performs sequence-level optimization to improve stability and efficiency, while coinciding with GRPO at the first iteration.
We identify two key issues in GRPO: token-level penalization and policy collapse, for which we show that KL divergence provides limited protection, while entropy-based analysis offers clearer corrective signals cui2025entropymechanismreinforcementlearning. These insights motivate GTPO, which improves stability and performance during training and evaluation.
@@ 3 Preliminaries
During training, given an input prompt $q$, the sampling LLM $\pi_{\theta_{\text{old}}}(\cdot\mid q)$ generates a group of $G$ completions $\{o_{i}\}_{i=1}^{G}$, where each completion consists of tokens denoted by $o_{i,t}$. Each completion receives a scalar reward $R_{i}$, combining accuracy and formatting verifiable scores. These rewards are then normalized into advantages $\hat{A}_{i}=(R_{i}-\bar{R})/{\mathrm{std}(R)}$. where $\bar{R}$ and $\mathrm{std}(R)$ denote the mean and standard deviation of the $R_{i}$ values within the group $G$. The GRPO objective is:
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{o_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\bar{\mathcal{C}}_{i}-\beta\cdot D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})\right],$ (1)
where $D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$ refers to KL divergence scaled by $\beta$ and $\bar{\mathcal{C}}_{i}$ is the average over tokens:
$\bar{\mathcal{C}}_{i}=\sum_{t=1}^{|o_{i}|}\frac{\min\Big(r_{i,t}(\theta)\,\hat{A}_{i},\;\operatorname{clip}\big(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\big)\,\hat{A}_{i}\Big)}{|o_{i}|},$ (2)
where the importance sampling ratio is defined as: $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid s_{i,t})},$ and the state at step $t$ is $s_{i,t}=(q,o_{i,<t})$, with $o_{i,<t}$ denoting the sequence of tokens generated before step $t$ for completion $i$.
The policy $\pi_{\theta}$ is obtained from logits $f_{\theta}$ via $\pi_{\theta}(o_{i,t}\mid s_{i,t})=\mathrm{softmax}(f_{\theta})\in\mathbb{R}^{V}$, where $V$ is the vocabulary size. The KL term in Eq. (1) penalizes deviations from a reference policy $\pi_{\mathrm{ref}}$, scaled by $\beta$. For clarity, in the gradient expression we omit the clipping and keep only the dependence on the ratio $r_{i,t}$ (full derivation in App. 8.1). Under the standard policy-gradient approximation,
@@
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Bigg(\frac{\hat{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}g_{i,t}$
$\;-\;\beta\cdot\nabla_{\theta}\mathrm{D}_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Bigg)\Bigg]$ (3)
Where the token-level contribution $g_{i,t}$ depends only on the ratio and on the policy gradient:
$g_{i,t}:=r_{i,t}(\theta)\,\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid s_{i,t})$ (4)
Writing $g_{i,t}$ in terms of logits, let $j$ be the index of $o_{i,t}$, with probability $\pi_{\theta}^{j}$; then
$g_{i,t}=r_{i,t}(\theta)\left[\nabla_{\theta}f_{\theta}^{j}(1-\pi_{\theta}^{j})-\sum_{k\neq j}\pi_{\theta}^{k}\,\nabla_{\theta}f_{\theta}^{k}\right]$ (5)
A positive normalized advantage $\hat{A}_{i}>0$ increases the logit of the selected token (scaled by $r_{i,t}$) while decreasing those of its alternatives; negative advantages have the opposite effect.
@@ 4 GRPO Issues
We highlight two major limitations of GRPO: token-level penalization and policy collapse.
@@ 4.1 Token-level Penalization
In this section we show that GRPO tends to have negative effects on the updates of tokens shared across multiple completions in the same group even though they can be essential for reward maximization.
@@
Prefix Tokens Penalization. Given a prompt $q$, the model generates a set of completions $\{o_{i}\}_{i=1}^{G}$, which may begin with the same sequence of tokens and then diverge at a specific point, a token that acts as a crossroads. This behavior is typical in instruction-tuned models, where early tokens tend to be similar across completions first-token-similar. However, not all completions in the group $G$ necessarily share the same prefix. In practice, we often observe multiple distinct prefixes, each shared by a subset of the completions. To model this, we partition the $G$ completions into $K$ disjoint groups $\mathcal{G}_{1},\ldots,\mathcal{G}_{K}$ so that all completions in a group $\mathcal{G}_{k}$ have a common prefix $S_{\text{pfx}}^{(k)}$. If all $G$ completions share the same prefix, then $K=1$. Tokens in $S_{\text{pfx}}^{(k)}$ are affected by the combined advantages of all completions in the group, while other tokens only depend on the advantage of their own completions. Under this structure, we rewrite the GRPO gradient (Eq. 3) by summing the contributions of all groups, where each group contributes a prefix term and a set of individual terms. The gradient, excluding the KL term, becomes:
@@
$\begin{split}\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{k=1}^{K}\Bigg[\sum_{i\in\mathcal{G}_{k}}\Bigg(\frac{\hat{\mathcal{A}}_{i}}{|{o}_{i}|}\sum_{j\notin S_{\text{pfx}}^{(k)}}g_{i,j}\Bigg)\\ &+\smash{\underbrace{\sum_{i\in\mathcal{G}_{k}}\Bigg(\frac{\hat{\mathcal{A}}_{i}}{|{o}_{i}|}\sum_{t\in S_{\text{pfx}}^{(k)}}g_{i,t}}_{\Delta_{\text{pfx}}^{(k)}}}\Bigg)\qquad\Bigg]\Bigg]\end{split}$ (6)
@@
The term $\Delta_{\text{pfx}}^{(k)}$ represents the gradient update associated with the shared prefix of group $\mathcal{G}_{k}$. If a completion does not share any prefix with others, then $\mathcal{G}_{k}$ contains only that single completion, resulting in $S_{\text{pfx}}^{(k)}=\emptyset$ and $\Delta_{\text{pfx}}^{(k)}=0$. Since the log-probability terms $g_{i,t}$ are identical and so constant across completions in the group $S_{\text{pfx}}^{(k)}$, the gradient component $\Delta_{\text{pfx}}$ is primarily driven by the sum of normalized advantages $\sum_{i\in\mathcal{G}_{k}}\hat{\mathcal{A}}_{i}/|{o}_{i}|$. This implies that the update to a prefix mainly depends on the relative balance of advantages across all completions in group $k$. For instance, the sum can be negative, when correct completions (with positive advantages) are generally longer than incorrect ones. In such cases, the denominator $|{o}_{i}|$ for the correct completions becomes larger, reducing their contribution to the overall gradient. As a result, the model may be penalized for generating desirable and beneficial prefix tokens. This introduces a bias against shared initial tokens, such as formatting tokens or reasoning tags (<reasoning>), essential for the response structure and correctness.
@@
Dependencies From Completion Advantage. GRPO can also induce undesirable penalization on tokens that appear after the prefix. Certain tokens, in particular formatting ones, may occur in multiple completions at varying positions and within different contexts, some correct, others incorrect, making them more susceptible to inconsistent and potentially harmful updates. This induces potential issues in the update. For instance, if a token $\tau$ frequently appears in completions with negative advantage, its probability may be reduced, even if the token is syntactically correct and required. Additionally, further issues can arise when a completion only partially follows the expected format. For example, if a completion closes a first part correctly with a formatting token $\tau$ (e.g. </reasoning>) but then omits <answer>, $\tau$ may receive a low or even negative reward. This, in turn, penalizes </reasoning>, despite it being a correct token in the completion. This occurs because the update for each token primarily depends on the advantage assigned to the entire completions in which it appears, without directly considering whether the token’s individual contribution was beneficial. This phenomenon is particularly impactful in the case of formatting tokens, as illustrated in the previous examples, and in the case of shared final tokens that form a common suffix among completions.